OpenStudy (anonymous):

If you weigh 400 pounds, lose 300 pounds, gain 1000 more pounds times 3, and divide it by 3.2, how much do you weigh on the moon?

OpenStudy (anonymous):

wouldn't that be 1031.somethin

OpenStudy (anonymous):

31.1 ponds

OpenStudy (anonymous):

but on the moon you weigh less so handy seems right

OpenStudy (anonymous):

400-300=100x3=300/3.2=187.5 then some fancy crap about gravity that i googled

OpenStudy (anonymous):

The moon is 1/4 the size of Earth, so the moon's gravity is much less than the earth's gravity, 83.3% (or 5/6) less to be exact.

OpenStudy (anonymous):

gotta love google right xD

OpenStudy (anonymous):

Heck yeah just Google that no mater what it is!!

OpenStudy (anonymous):

Okay thanks guys! ♥